15. Accounts receivable Effective credit management involves establishing credit standards for extending credit to customers, determining the companys terms of credit, and setting up procedures for invoicing and collecting past-due accounts. The following statement refers to a credit management policy. Select the best term to complete the sentence. How a company handles its credit accounts, indluding methods of invoicing and collecting past-due accounts, is indicated by the companys Consider the case of Ziff Corp.: Ziff Corp. has a very attractive credit policy, and none of its customers pays in cash when the firm makes a sale. Ziff Corp. sells to its customers on credit terms of 1/10, net 30 If a customer bought $150,000 worth of goods and paid the firm cash eight days after the sale, how much cash would Ziff Corp. get from the customer? O $148,500 O $123,750 O $127,500 O $157,50o If the customer paid off the account after 15 days, Ziff Corp. would receive Approximately 30% of Ziff Corp.s customers take advantage of the discount and pay on the 10th day. The remaining 70% take an average of 35 days to pay off their accounts. what is Ziff Corp.s days sales outstanding (DSO), or the average collection period? O 31.6 days 23.4 days O 30.3 days O 27.5 days
Answer #1

Is indicated by the company’s DSO (days sales outstanding)/Average Collection Period

Credit Term 1/10 net 30 means discount of 1 percent if payment made within 10 days, otherwise payment has to be made within 30 days with no discount

Cash collected = 150,000*99% = $148,500

If paid after 15 days, no discount will be given , amount will be $150,000

Average Collection Period = 0.3*10 + 0.7*35 = 27.5 days